OpenStudy (anonymous):

expand and simplify where appropriate (x-8)^2

OpenStudy (anonymous):

HELPPPP

OpenStudy (anonymous):

(x-8)(x-8) = x ^2 -8x+ (-8x) (-)-8^2 = x^2 - 16x + 64

OpenStudy (anonymous):

is that what you need to do O. o?

OpenStudy (anonymous):

yeah, i did it again and got that but i didnt know if it was right.. thanks heaps

OpenStudy (anonymous):

:D you are welcome~

OpenStudy (anonymous):

i have another question. Do you know how to do parabolas??

OpenStudy (anonymous):

um, i will try :D

OpenStudy (anonymous):

ok, because i keep doing something wrong and i was just wondering if you could do the equation so i could see where i went wrong.. if thats not asking too much??

OpenStudy (anonymous):

yeah, sure, i will do my best XD

OpenStudy (anonymous):

ok, thanks. heres the question..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Over a period of 8hours, the temperature of a room follows the relationship T=h^2-6h+15 where T is the temperature in degrees Celcius, h hours after commencing the experiment a) Sketch the graph of this quadratic relationship b) What is the initial temperature? c) After 2 hours, is the temperature of the room increasing or decreasing? d) After 6 hours is the temperature of the room increasing or decreasing? e) Find the minimum temperature f) When did the minimum temperature occur? Sorry, it's alot :(

OpenStudy (anonymous):

wait XD

OpenStudy (anonymous):

a) y intercept : h=0 T=y y= 0^2 - 6(0) + 15 = 15  i think the initial temperature will be the y intercept (0, 15) for the hours(x/ h)can not be negative c) T = (2)^2 - 6(2) + 15 (sub in or look at the graph) = 4 - 12 +15 = 7 (which mean decrease) d) same sub in or look at the graph T = (6)^2 - 6(6) + 15 = 36 - 36 + 15 = 15 (which mean it stay the same ( so it goes decrease at h = 7 and slowly go back up and h=6 will be back to the start temperature ) e) the minimum is the turning point which is T = x^2 - 6h + 15 = (x^2-6x+9) + 6 = (x - 3)^2 + 6 TP: (3, 6)